# NVINS: Robust Visual Inertial Navigation Fused with NeRF-augmented Camera Pose Regressor and Uncertainty Quantification

### v1

@@ Abstract
In recent years, Neural Radiance Fields (NeRF) have emerged as a powerful tool for 3D reconstruction and novel view synthesis. However, the computational cost of NeRF rendering and degradation in quality due to the presence of artifacts pose significant challenges for its application in real-time and robust robotic tasks, especially on embedded systems. This paper introduces a novel framework that integrates NeRF-derived localization information with Visual-Inertial Odometry (VIO) to provide a robust solution for robotic navigation in a real-time. By training an absolute pose regression network with augmented image data rendered from a NeRF and quantifying its uncertainty, our approach effectively counters positional drift and enhances system reliability. We also establish a mathematically sound foundation for combining visual inertial navigation with camera localization neural networks, considering uncertainty under a Bayesian framework. Experimental validation in the photorealistic simulation environment demonstrates significant improvements in accuracy compared to a conventional VIO approach.
@@ II Related Work
Uncertainty in camera pose regression is captured through covariance matrix of the pose, distinguishing between aleatoric and epistemic uncertainty [26]. Aleatoric uncertainty refers to the irreducible uncertainty associated with the noise inherent in the measurements. Epistemic uncertainty is the reducible uncertainty stemming from a lack of knowledge beyond the training dataset. Kendall et al. [21] quantify epistemic uncertainty using MC dropout to estimate errors caused by out-of-distribution inputs. To combine uncertainty estimated by the pose regressor with additional sensor measurements, conventional methods such as the Kalman filter [27] or visual odometry (VO) [28] have been used. Although epistemic and aleatoric uncertainty have been simultaneously estimated in prior work [28], we are not aware of any system which thoroughly utilizes this type of 6-DoF uncertainty estimate in a pose graph optimization framework.
@@ III Preliminaries
Under the BNN framework, the neural network’s weights, $\mathbf{w}$, are not considered as fixed values. Instead, they initially follow a prior probabilistic distribution, $P(\mathbf{w})$. For a given training dataset $\mathcal{D}=\{\mathcal{X},\mathcal{Y}\}$, where $\mathcal{X}$ represents the input dataset and $\mathcal{Y}$ the label dataset, these weights conform to a posterior distribution, $P(\mathbf{w}|\mathcal{D})$.
Deep Ensemble [17, 30]: In deep ensemble, a finite set of weights, $\mathcal{W}_{p}=\{\mathbf{w}_{i}^{p}\}$, is sampled from the prior distribution of weights, $P(\mathbf{w})$. Each sampled weight, $\mathbf{w}_{i}^{p}$, is optimized to find a maximum a posteriori (MAP) estimate $\mathbf{w}_{i}^{*}=\operatorname*{argmax}_{\mathbf{w}}P(\mathbf{w}|\mathcal{D})=\operatorname*{argmin}_{\mathbf{w}}(-\log P(\mathcal{Y}|\mathbf{w},\mathcal{X})P(\mathbf{w}))$. Then, the posterior distribution $P(\mathbf{w}|\mathcal{D})$ is approximated as
@@
where the two components, $\mathcal{L}(\mathbf{w})$ and $\mathcal{L}_{reg}(\mathbf{w})$, correspond to the negative log likelihood of $P(\mathcal{Y}|\mathbf{w},\mathcal{X})$ and $P(\mathbf{w})$, respectively.
Monte Carlo (MC) dropout [16]: MC dropout is a variational inferenece method which approximates the true posterior distribution over weights $P(\mathbf{w}|\mathcal{D})$ as $q_{\theta^{*}}(\mathbf{w})$, a neural network with model parameters $\theta^{*}$. At test time, weights are sampled from $q_{\theta^{*}}(\mathbf{w})$, i.e., $\hat{\mathbf{w}}=\{\hat{\mathbf{W}}_{i}\}_{i=1}^{L}\sim q_{\theta^{*}}(\mathbf{w})$. The weight matrix of $i$-th layer, $\hat{\mathbf{W}}_{i}$, is dropped out with a probability $1-p_{i}$. That is, the weights at each layer $i$ follow the Bernoulli distribution:
@@ IV-A Training NeRF and Image Rendering
We use a NeRF reconstruction as a form of data augmentation, generating a large dataset $\mathcal{D}_{A}=\{\mathcal{I}_{A},\mathcal{T}_{A}\}$ from a relatively small dataset $\mathcal{D}_{O}=\{\mathcal{I}_{O},\mathcal{T}_{O}\}$. A neural network $[\mathbf{c},\sigma]=f_{\phi}(\mathbf{x},\mathbf{d})$ parameterized by $\phi$ is trained with $\mathcal{D}_{A}$. The network takes two inputs: $\mathbf{x}\in\mathbb{R}^{3}$ and $\mathbf{d}\in\mathbb{R}^{2}$ representing the position and the direction of the ray for rendering, respectively. The outputs, $\mathbf{c}$ and $\sigma$, represent the view-dependent color and view-independent volume density of the sampled point of the ray. An intensity of one pixel is obtained through rendering a ray. The loss function of the neural network $f_{\phi}(\mathbf{x},\mathbf{d})$ is designed to minimize the gap between the pixel intensity of the camera image and the estimated pixel intensity obtained through rendering process. For more information about NeRF, we refer to [1].
@@ IV-B Camera Pose Regression with Uncertainty Quantification
$I_{n}\in\mathbb{R}^{W\times H\times 3}$ is the input of the function $f_{\mathbf{w}}$. $f_{\mathbf{w}}$ utilizes a CNN to process high-dimensional image data. Notably, pixel coordinates are added as extra channels before each convolutional layer to improve CNN performance on coordinate transformation-related tasks [31]. The vectorized pose representation of $T_{n}$, $\hat{\mathbf{t}}_{n}=[\hat{\mathbf{p}}_{n}^{T}\;\hat{\mathbf{r}}_{n}^{T}]^{T}\in\mathbb{R}^{9}$, consists of a 3D position and a 6D representation of rotation. This choice is made to avoid the high errors associated with discontinuous representations such as axis-angle or quaternion [32]. We approximate the covariance of the position $\mathbf{p}_{n}$ as a diagonal matrix
$S_{n,p}^{a}=\text{diag}(\boldsymbol{\sigma}_{n,p}^{a}).$ (6)
@@
$P(\mathbf{t}_{n}|I_{n},\mathbf{w})=\frac{\exp\big{(}-\frac{1}{2}\|\mathbf{p}_{n}-\hat{\mathbf{p}}_{n}\|_{S^{a}_{n}}^{2}+\kappa\text{tr}(\hat{R}_{n}^{T}R_{n})\big{)}}{c(\kappa)\sqrt{(2\pi)^{3}\det{(S^{a}_{n})}}}$ (7)
@@
We leverage the BNN to model an epistemic uncertainty, which arises when the model encounters inputs that deviate from the the training dataset $\mathcal{D}_{A}$. Unlike conventional neural networks, in a BNN, the weight $\mathbf{w}$ does not assume a fixed value but follows a posterior weight distribution $P(\mathbf{w}|\mathcal{D}_{A})$ given the dataset $\mathcal{D}_{A}$, as discussed in Sec. III. To approximate this posterior weight distribution, we adopt deep ensemble [17] and MC dropout [16] techniques due to their simplicity and scalability compared to other approaches [34].
The loss function for the weight $\mathbf{w}$ is formulated as shown in (2). Assuming independence between each data pair $(I_{n},\mathbf{t}_{n})$, the loss function $\mathcal{L}(\mathbf{w})$ corresponding to the negative log likelihood of $P(\mathcal{T}_{A}|\mathcal{I}_{A},\mathbf{w})$ is represented as
@@
For numerical stability, the covariance output of the network $\boldsymbol{\sigma}_{n}^{a}$ is replaced with the log vector of the covariance output, $\mathbf{s}_{n,p}^{a}$, such that $\sigma_{n,p,m}^{a}=\exp(s_{n,p,m}^{a})$. Thus, equation (8) is rewritten as:
$\begin{split}&\mathcal{L}(\mathbf{w})=\frac{1}{N}\sum_{n=1}^{N}\frac{1}{2}\Big{(}\kappa\|R_{n}-\hat{R}_{n}\|^{2}_{F}+\\ &\sum_{m=1}^{3}\big{(}e^{-s^{a}_{n,p,m}}\|p_{n,i}-\hat{p}_{n,i}\|^{2}_{2}+s^{a}_{n,p,m}(I_{n})\big{)}\Big{)}\end{split}$ (9)
Both the estimated mean, $\hat{\mathbf{t}}_{n}$, and log of variance, $\mathbf{s}_{n,p}^{a}$, are parameterized by $\mathbf{w}$. Additionally, the regularization loss, $\mathcal{L}_{reg}(\mathbf{w})$, assumes the prior distribution $P(\mathbf{w})$ is Gaussian with zero mean, formulated as $\mathcal{L}_{reg}(\mathbf{w})=\lambda\|\mathbf{w}\|_{2}^{2}$ with a positive constant $\lambda$.
@@
For practical implementation, a weight $\mathbf{w}_{i}$ is sampled from the distribution $q(\mathbf{w})$, approximating $P(\mathbf{w}|\mathcal{D}_{A})$ through either deep ensemble or MC dropout. This results in a set of weights $\mathcal{W}=\{\mathbf{w}_{i}\}$. The pose distribution $P(\mathbf{t}_{n}|\mathcal{D}_{A},I_{n})$ is parameterized as the average and covariance of the outputs from neural networks with the sampled weights, represented as $[(\mathbf{t}_{n}^{i})^{T}\;(\boldsymbol{\sigma}_{n,p}^{a,i})^{T}]^{T}=f_{\mathbf{w}_{i}}(I_{n}),\ \forall\mathbf{w}_{i}\in\mathcal{W}$. The averages of the position and the orientation are estimated as:
@@
The estimated covariance, $\hat{\Sigma}_{n,\{p,r\}}=\hat{\Sigma}_{n,\{p,r\}}^{a}+\hat{\Sigma}_{n,\{p,r\}}^{e}$, of the pose distribution (10) is estimated as the sum of the estimated aleatoric uncertainty, $\hat{\Sigma}_{n,\{p,r\}}^{a}$, and the estimated epistemic uncertainty, $\hat{\Sigma}_{n,\{p,r\}}^{e}$ reflecting the distribution (7) with $\{p,\,r\}$ representing position and rotation [26]. Assuming the distribution of $\|R_{n}-\bar{R}_{n}\|_{F}$ can be approximated by a zero-mean Gaussian distribution, the uncertainties can be represented as follows:
$\begin{gathered}\hat{\Sigma}_{n,p}^{a}=\frac{\sum_{w_{i}\in\mathcal{W}}S_{n}^{a,i}}{|\mathcal{W}|-1},\hat{\Sigma}_{n,p}^{e}=\frac{\sum_{w_{i}\in\mathcal{W}}(\mathbf{p}_{n}^{i}-\bar{\mathbf{p}}_{n})(\mathbf{p}_{n}^{i}-\bar{\mathbf{p}}_{n})^{T}}{|\mathcal{W}|-1}\\ \hat{\Sigma}_{n,r}^{a}=\kappa^{-1},\hat{\Sigma}_{n,r}^{e}=\frac{\sum_{w_{i}\in\mathcal{W}}\|R_{n}^{i}-\bar{R}_{n}\|_{F}^{2}}{|\mathcal{W}|-1}.\end{gathered}$ (12)
@@ V VIO Integration with Camera Pose Regressor
Moreover, we selectively exclude the neural networks output identified as outliers from the optimization, ensuring a more robust MAP estimation using the uncertainty of the neural network.
@@ V-A Factor Graph Optimization
Assuming zero-mean Gaussian noise for all measurements and that all measurements are independent of each other except the relation between the offline data $\mathcal{D}_{A}$ and camera images $\mathcal{I}_{o}$, the MAP estimate is given by
$\begin{split}&\mathcal{X}^{*}=\operatorname*{argmax}_{\mathcal{X}}P(\mathcal{X}|\mathcal{D}_{A},\mathcal{Z})\\ &=\operatorname*{argmax}_{\mathcal{X}}P(\mathcal{X}|\mathcal{D}_{A},\mathcal{I}_{o})P(\mathcal{Z}_{B},\mathcal{Z}_{C}|\mathcal{X},\mathcal{D}_{A},\mathcal{I}_{o})\\ &=\operatorname*{argmax}_{\mathcal{X}}P(\mathcal{X}|\mathcal{D}_{A},\mathcal{I}_{o})P(\mathcal{Z}_{B}|\mathcal{X})P(\mathcal{Z}_{C}|\mathcal{X})\\ &=\operatorname*{argmax}_{\mathcal{X}}\prod_{n\in\mathcal{N}}P(T_{n}|\mathcal{D}_{A},I_{n})\prod_{b\in\mathcal{B}}P(\mathbf{z}_{b_{k+1}}^{b_{k}}|\mathcal{X})\prod_{(i,j)\in\mathcal{C}}P(\mathbf{z}^{j}_{i}|\mathcal{X}).\end{split}$ (16)
@@
The nonlinear optimization problem (16) is solved by minimizing the sum of Mahalanobis distances
$\begin{split}\mathcal{X}^{*}&=\operatorname*{argmin}_{\mathcal{X}}\Big{\{}\sum_{n\in\mathcal{N}}\|\mathbf{r}_{\mathcal{N}}(I_{n},\mathcal{X})\|_{\hat{\Sigma}_{n}}^{2}\\ &+\sum_{k\in\mathcal{B}}\|\mathbf{r}_{\mathcal{B}}(\mathbf{z}^{b_{k}}_{b_{k+1}},\mathcal{X})\|_{\Sigma^{b_{k}}_{b_{k+1}}}^{2}+\sum_{(i,j)\in\mathcal{C}}\|\mathbf{r}_{\mathcal{C}}(\mathbf{z}^{j}_{i},\mathcal{X})\|_{\Sigma^{j}_{l}}^{2}\Big{\}}\end{split}$ (17)
@@
$\mathbf{r}_{\mathcal{N}}(I_{n},\mathcal{X})=[(\mathbf{p}_{n}-\bar{\mathbf{p}}_{n})^{T}\,\|R_{n}-\bar{R}_{n}\|_{F}]^{T}.$ (18)

### v2

@@ Abstract
In recent years, Neural Radiance Fields (NeRF) have emerged as a powerful tool for 3D reconstruction and novel view synthesis. However, the computational cost of NeRF rendering and degradation in quality due to the presence of artifacts pose significant challenges for its application in real-time and robust robotic tasks, especially on embedded systems. This paper introduces a novel framework that integrates NeRF-derived localization information with Visual-Inertial Odometry (VIO) to provide a robust solution for real-time robotic navigation. By training an absolute pose regression network with augmented image data rendered from a NeRF and quantifying its uncertainty, our approach effectively counters positional drift and enhances system reliability. We also establish a mathematically sound foundation for combining visual inertial navigation with camera localization neural networks, considering uncertainty under a Bayesian framework. Experimental validation in a photorealistic simulation environment demonstrates significant improvements in accuracy compared to a conventional VIO approach.
@@ II Related Work
Uncertainty in camera pose regression is captured through the covariance matrix of the pose, distinguishing between aleatoric and epistemic uncertainty [26]. Aleatoric uncertainty refers to the irreducible uncertainty associated with the noise inherent in the measurements. Epistemic uncertainty is the reducible uncertainty stemming from a lack of knowledge beyond the training dataset. Kendall et al. [21] quantify epistemic uncertainty using MC dropout to estimate errors caused by out-of-distribution inputs. To combine uncertainty estimated by the pose regressor with additional sensor measurements, conventional methods such as the Kalman filter [27] or visual odometry (VO) [28] have been used. Although epistemic and aleatoric uncertainty have been simultaneously estimated in prior work [28], we are not aware of any system that thoroughly utilizes this type of 6-DoF uncertainty estimate in a pose graph optimization framework.
@@ III Preliminaries
Under the BNN framework, the neural network’s weights, $\mathbf{w}$, are not considered fixed values. Instead, they initially follow a prior probabilistic distribution, $P(\mathbf{w})$. For a given training dataset $\mathcal{D}=\{\mathcal{X},\mathcal{Y}\}$, where $\mathcal{X}$ represents the input features and $\mathcal{Y}$ the target variable, these weights conform to a posterior distribution, $P(\mathbf{w}|\mathcal{D})$.
Deep Ensemble [17, 30]: In deep ensemble, a finite set of weights, $\mathcal{W}_{p}=\{\mathbf{w}_{i}^{p}\}$, is sampled from the prior distribution of weights, $P(\mathbf{w})$. Each sampled weight, $\mathbf{w}_{i}^{p}$, is optimized to find a maximum a posteriori (MAP) estimate $\mathbf{w}_{i}^{*}=\operatorname*{argmax}_{\mathbf{w}}P(\mathbf{w}|\mathcal{D})=\operatorname*{argmin}_{\mathbf{w}}(-\log P(\mathcal{Y}|\mathbf{w},\mathcal{X})P(\mathbf{w}))$ where $P(\mathcal{Y}|\mathbf{w},\mathcal{X})$ is a likelihood of the label data $\mathcal{Y}$ given the input data $\mathcal{X}$ and weights $\mathbf{w}$. Then, the posterior distribution $P(\mathbf{w}|\mathcal{D})$ is approximated as
@@
where the two components, $\mathcal{L}(\mathbf{w})$ and $\mathcal{L}_{reg}(\mathbf{w})$, correspond to the negative log-likelihood of $P(\mathcal{Y}|\mathbf{w},\mathcal{X})$ and $P(\mathbf{w})$, respectively.
Monte Carlo (MC) dropout [16]: MC dropout is a variational inference method which approximates the true posterior distribution over weights $P(\mathbf{w}|\mathcal{D})$ as $q_{\theta^{*}}(\mathbf{w})$, a neural network with model parameters $\theta^{*}$. At test time, weights are sampled from $q_{\theta^{*}}(\mathbf{w})$, i.e., $\hat{\mathbf{w}}=\{\hat{\mathbf{W}}_{i}\}_{i=1}^{L}\sim q_{\theta^{*}}(\mathbf{w})$. The weight matrix of $i$-th layer, $\hat{\mathbf{W}}_{i}$, is dropped out with a probability $1-p_{i}$. That is, the weights at each layer $i$ follow the Bernoulli distribution:
@@ IV-A Training NeRF and Image Rendering
We use a NeRF reconstruction as a form of data augmentation, generating a large dataset $\mathcal{D}_{A}=\{\mathcal{I}_{A},\mathcal{T}_{A}\}$ from a relatively small dataset $\mathcal{D}_{O}=\{\mathcal{I}_{O},\mathcal{T}_{O}\}$. Even if the original poses $\mathcal{T}_{O}$ are not accurate, they can be refined using bundle adjustment techniques such as COLMAP [31]. A neural network $[\mathbf{c},\sigma]=f_{\phi}(\mathbf{x},\mathbf{d})$ parameterized by $\phi$ is trained with $\mathcal{D}_{O}$. The network takes two inputs: $\mathbf{x}\in\mathbb{R}^{3}$ and $\mathbf{d}\in\mathbb{R}^{2}$ representing the position and the direction of the ray for rendering, respectively. The outputs, $\mathbf{c}$ and $\sigma$, represent the view-dependent color and view-independent volume density of the sampled point of the ray. An intensity of one pixel is obtained through rendering a ray. The loss function of the neural network $f_{\phi}(\mathbf{x},\mathbf{d})$ is designed to minimize the gap between the pixel intensity of the camera image and the estimated pixel intensity obtained through rendering process. For more information about NeRF, we refer to [1].
@@ IV-B Camera Pose Regression with Uncertainty Quantification
$I_{n}\in\mathbb{R}^{W\times H\times 3}$ is the input of the function $f_{\mathbf{w}}$. $\hat{\mathbf{t}}_{n}$ and $\boldsymbol{\sigma}_{n,p}^{a}$ are the vectorized representation of the pose and aleatoric uncertainty of the position, respectively. $f_{\mathbf{w}}$ utilizes a CNN to process high-dimensional image data. Notably, pixel coordinates are added as extra channels before each convolutional layer to improve CNN performance on coordinate transformation-related tasks [32]. The vectorized pose representation, $\hat{\mathbf{t}}_{n}=[\hat{\mathbf{p}}_{n}^{T},\hat{\mathbf{r}}_{n}^{T}]^{T}\in\mathbb{R}^{9}$, consists of a 3D position and a 6D representation of rotation. This choice avoid the high errors associated with discontinuous representations such as axis-angle or quaternion [33]. We approximate the covariance of the position $\mathbf{p}_{n}$ as a diagonal matrix, i.e. $S_{n,p}^{a}=\text{diag}(\boldsymbol{\sigma}_{n,p}^{a})$.
@@
$P(\mathbf{t}_{n}|I_{n},\mathbf{w})=\frac{\exp\big{(}-\frac{1}{2}\|\mathbf{p}_{n}-\hat{\mathbf{p}}_{n}\|_{S^{a}_{n}}^{2}+\kappa\text{tr}(\hat{R}_{n}^{T}R_{n})\big{)}}{c(\kappa)\sqrt{(2\pi)^{3}\det{(S^{a}_{n,p})}}}$ (6)
@@
We leverage the BNN to model an epistemic uncertainty, which arises when the model encounters inputs that deviate from the training dataset $\mathcal{D}_{A}$. As shown in Fig. 2, rendered images $\mathcal{I_{A}}$ cannot be perfectly identical to the original. When the model takes the images in the real environment, the gap between the rendered images and the original images is modeled as epistemic uncertainty. Unlike conventional neural networks, in a BNN, the weight $\mathbf{w}$ does not assume a fixed value but follows a posterior weight distribution $P(\mathbf{w}|\mathcal{D}_{A})$ given the dataset $\mathcal{D}_{A}$, as discussed in Sec. III. To approximate the posterior weight distribution, we adopt deep ensemble [17] and MC dropout [16] due to their simplicity and scalability compared to other approaches [35].
The loss function for the weight $\mathbf{w}$ is formulated as shown in (2). Assuming independence between each data pair $(I_{n},\mathbf{t}_{n})$, the loss function $\mathcal{L}(\mathbf{w})$ corresponding to the negative log-likelihood of $P(\mathcal{T}_{A}|\mathcal{I}_{A},\mathbf{w})$ is represented as
@@
For numerical stability, the covariance output of the network $\boldsymbol{\sigma}_{n}^{a}$ is replaced with the log vector of the covariance output, $\mathbf{s}_{n,p}^{a}$, such that $\sigma_{n,p,m}^{a}=\exp(s_{n,p,m}^{a})$. Both the estimated mean, $\hat{\mathbf{t}}_{n}$, and log of variance, $\mathbf{s}_{n,p}^{a}$, are parameterized by $\mathbf{w}$. Additionally, the regularization loss, $\mathcal{L}_{reg}(\mathbf{w})$, assumes the prior distribution $P(\mathbf{w})$ is Gaussian with zero mean, formulated as $\mathcal{L}_{reg}(\mathbf{w})=\lambda\|\mathbf{w}\|_{2}^{2}$ with a positive constant $\lambda$.
@@
For practical implementation, a weight $\mathbf{w}_{i}$ is sampled from the distribution $q(\mathbf{w})$, approximating $P(\mathbf{w}|\mathcal{D}_{A})$ through either deep ensemble or MC dropout. This results in a set of weights $\mathcal{W}=\{\mathbf{w}_{i}\}$. The pose distribution $P(\mathbf{t}_{n}|\mathcal{D}_{A},I_{n})$ is parameterized as the average and covariance of the outputs from neural networks with the sampled weights, represented as $f_{\mathbf{w}_{i}}(I_{n})=[(\mathbf{t}_{n}^{i})^{T}\;(\boldsymbol{\sigma}_{n,p}^{a,i})^{T}]^{T},\ \forall\mathbf{w}_{i}\in\mathcal{W}$. The averages of the position and the orientation are estimated as:
@@
The covariance, $\hat{\Sigma}_{n,\{p,r\}}=\hat{\Sigma}_{n,\{p,r\}}^{a}+\hat{\Sigma}_{n,\{p,r\}}^{e}$, of the pose distribution (8) is estimated as the sum of the aleatoric uncertainty, $\hat{\Sigma}_{n,\{p,r\}}^{a}$, and the epistemic uncertainty, $\hat{\Sigma}_{n,\{p,r\}}^{e}$ with $\{p,\,r\}$ representing position and rotation [26]. Assuming the distribution of $\|R_{n}-\bar{R}_{n}\|_{F}$ can be approximated by a zero-mean Gaussian distribution, the uncertainties can be represented as follows:
$\begin{gathered}\hat{\Sigma}_{n,p}^{a}=\frac{\sum_{w_{i}\in\mathcal{W}}S_{n}^{a,i}}{|\mathcal{W}|-1},\hat{\Sigma}_{n,p}^{e}=\frac{\sum_{w_{i}\in\mathcal{W}}\tilde{\mathbf{p}}_{n}^{i}(\tilde{\mathbf{p}}_{n}^{i})^{T}}{|\mathcal{W}|-1}\\ \hat{\Sigma}_{n,r}^{a}=\kappa^{-1},\hat{\Sigma}_{n,r}^{e}=\frac{\sum_{w_{i}\in\mathcal{W}}\|R_{n}^{i}-\bar{R}_{n}\|_{F}^{2}}{|\mathcal{W}|-1}\end{gathered}$ (10)
@@ V VIO Integration with Camera Pose Regressor
Moreover, we selectively exclude the neural network output identified as outliers from the optimization, ensuring a more robust MAP estimation using the uncertainty of the neural network.
@@ V-A Factor Graph Optimization
Assuming zero-mean Gaussian noise for all measurements and that all measurements are independent of each other except the relation between the offline data $\mathcal{D}_{A}$ and camera images $\mathcal{I}_{o}$, the MAP estimate (11) is given by
$\begin{split}\mathcal{X}^{*}=\operatorname*{argmax}_{\mathcal{X}}&P(\mathcal{X}|\mathcal{D}_{A},\mathcal{I}_{o})P(\mathcal{Z}_{B},\mathcal{Z}_{C}|\mathcal{X},\mathcal{D}_{A},\mathcal{I}_{o})\\ =\operatorname*{argmax}_{\mathcal{X}}&P(\mathcal{X}|\mathcal{D}_{A},\mathcal{I}_{o})P(\mathcal{Z}_{B}|\mathcal{X})P(\mathcal{Z}_{C}|\mathcal{X})\\ =\operatorname*{argmax}_{\mathcal{X}}&\Big{\{}\prod_{n\in\mathcal{N}}P(T_{n}|\mathcal{D}_{A},I_{n})\\ &\prod_{b\in\mathcal{B}}P(\mathbf{z}_{b_{k+1}}^{b_{k}}|\mathcal{X})\prod_{(i,j)\in\mathcal{C}}P(\mathbf{z}^{j}_{i}|\mathcal{X})\Big{\}}.\end{split}$ (14)
@@
The nonlinear optimization problem (14) is solved by minimizing the sum of Mahalanobis distances:
$\begin{split}&\mathcal{X}^{*}=\operatorname*{argmin}_{\mathcal{X}}\Big{\{}\sum_{n\in\mathcal{N}}\|\mathbf{r}_{\mathcal{N}}(I_{n},\mathcal{X})\|_{\hat{\Sigma}_{n}}^{2}\\ &+\sum_{k\in\mathcal{B}}\|\mathbf{r}_{\mathcal{B}}(\mathbf{z}^{b_{k}}_{b_{k+1}},\mathcal{X})\|_{\Sigma^{b_{k}}_{b_{k+1}}}^{2}+\sum_{(i,j)\in\mathcal{C}}\|\mathbf{r}_{\mathcal{C}}(\mathbf{z}^{j}_{i},\mathcal{X})\|_{\Sigma^{j}_{l}}^{2}\Big{\}}\end{split}$ (15)
@@
$\mathbf{r}_{\mathcal{N}}(I_{n},\mathcal{X})=[(\mathbf{p}_{n}-\bar{\mathbf{p}}_{n})^{T},\,\|R_{n}-\bar{R}_{n}\|_{F}]^{T}.$ (16)